Mark is 3 less than 6 times Pat's age. Mark is 35 years old. How old is Pat?
Answer by edjones(8007)

p=pat's age
6p-3=35
add 3 to each side: 6p=38
divide 6 into each side: p=6 1/3 y.o.
Check:
6*19/3-3=35
38-3=35 true
Ed